At the same store, Leigh bought 2 pairs of pants and 5 shirts for $61, and Jessica bought 3 pairs of pants and 2 shirts for $64. How much does 1 pair of pants cost?
Answer by addingup(3677)

2x+5y = 61 Multiply both sides time 3 to create a new equation
3x+2y = 64 Multiply both sides times 2 to create a new equation
Create the new equations and subtract one from the other, like this:
6x+15y = 183
-
6x+4y = 128
--------------
0x+11y = 55 Here 0x means 0, so eliminate it
11y = 55
y = 55/11 = 5 each shirt is 5
2x+5(5) = 61
2x+25 = 61
2x = 36
x = 36/2 = 18 each pair of pants is 18
----------------------------------
Check: Put these results in the original equations:
2x+5y = 61
2(18)+5(5) = 61
36+25 = 61 Correct
Next:
3x+2y = 64
3(18)+2(5) = 64
54+10 = 64 Also correct
:
John